OpenStudy (anonymous):

solve for g. b=1/2gw

OpenStudy (compassionate):

g = 2b/w

OpenStudy (anonymous):

multiply by 2 then divide by w tell me what will u get?

OpenStudy (lgbasallote):

\[b = \frac{1}{2gw}\] cross multiply \[2bgw = 1\] now divide both sides by 2 \[bgw = \frac 12\] divide both sides by b \[gw = \frac{1}{2b}\] divide both sides by w \[g = \frac{1}{2bw}\]